On June 4, Donaldson's Trucking accepts a $2600 note in settlement of a bill. The note is for 180 days at 8%
interest. Donaldson sells the note at a 11% discount rate on July 22. Find the proceeds.
1
Expert's answer
2018-12-06T11:09:10-0500

Interest receivable after 180 days (at maturity):

2,600*0.08*180/360=$104

Total receivable after 180 days:

2,600+104=$2,704

Discount at 11% on July 22:

2,704*0.11*132/360=$109

where 132 - the number of days remaining until maturity at the time of sale of the note:

180 – 48 (26 days in June and 22 in July) = 132

Therefore the discount fee will be more than interest receivable at maturity, the proceeds at the moment of selling will be less than face price of the note:

2,704-109=$2,595